# Counting Guidance for High Fidelity Text-to-Image Synthesis

### v1

@@ Introduction
The second row of Fig. 3 shows the effectiveness of our attention map guidance. We find that the attention map of “oranges” only focuses on oranges, and the attention map of “eggs” only focuses on eggs, resulting in a correctly synthesized output. Furthermore, we observe that high-fidelity object masks are generated from the corresponding attention maps.
Masked Counting Guidance
@@
$M^{b}_{i,j,k}=\begin{cases}1,&\text{if}\ i=\operatorname*{argmax}_{i}(M_{i,j,k})\\ 0,&\text{otherwise}\end{cases}$ (16)
@@
$L_{count,i}=|\frac{Count(\hat{x}_{0,i})-N_{i}}{N_{i}}|^{2}.$ (18)
@@
$\epsilon\leftarrow\epsilon+\sum_{i}s_{count,i}\sqrt{1-\bar{\alpha}_{t}}\nabla_{z_{t}}L_{count,i}$ (19)
@@
Experiments
We borrow the state-of-the-art text-to-image generation model Stable Diffusion v1.4 for our experiments. We use DDIM sampling with 50 steps and set the scale parameter for $L_{max}$ to $s_{max}=0.1$ by default. We create a modified dataset based on the object classes in Attend-and-Excite (Chefer et al. 2023) to evaluate and compare our approach with previous methods. Specifically, we remove the color category and add more animals and objects to focus on counting performance. We compare our method with Stable Diffusion (Rombach et al. 2022) and Attend-and-Excite.
@@
Fig. 4 shows a qualitative comparison for the single object type scenario. While Stable Diffusion and Attend-and-Excite fail to generate the same number of objects as specified in the prompt, our method generates the correct number. For the text prompt “four tomatoes on the table,” Stable Diffusion generates only three tomatoes without counting guidance. With counting guidance, the tomato at the bottom is successfully divided into two tomatoes, while the rest of the image is consistent with the original result. The text prompt “ten oranges on the table,” causes Stable Diffusion to only generate four oranges compared to our solution that creates the correct amount of ten. The big difference in count results in large gradients, making our result severely differ from the original.
Our method also works well for more complex categories, such as animals. Considering text prompt “three chicks on the road”, Stable Diffusion and Attend-and-Excite synthesize only two chicks, unlike our method which generates one additional chick while maintaining the other two chicks’ appearance. For the text prompt “five rabbits in the yard” Stable Diffusion and Attend-and-Excite generate only four rabbits, while our method generates one more rabbit but fails to maintain the other rabbits’ appearance. That is because of the difference between the background and the rabbit colors. It is hard to generate a white rabbit from a brown yard, so Stable Diffusion with counting guidance changes the overall structure and recreates five rabbits.
@@
Fig. 5 shows a qualitative comparison for multiple object types. For the given text prompt “three lemons and one bread on the table,” Stable Diffusion successfully creates one bread but fails to generate three lemons, while Attend-and-Excite fails in both cases. With masked counting guidance, our method correctly synthesizes three lemons and one bread. The result shows that the lemon at the bottom is divided into two lemons thanks to masked counting guidance while maintaining the bread’s shape.
@@
For the text prompt “two onions and two tomatoes on the table,” we find that Stable Diffusion suffers from the semantic information mixing problem and generates red onions instead of tomatoes. Due to our attention map guidance, our method creates real tomatoes. As Attend-and-Excite is also based on attention map optimization, it successfully generates real tomatoes but fails to generate the exact number of onions.
Limitations
@@
As our results show, our method aids in generating the exact number of each object, but we also found several limitations. First, it is often necessary to tune the scale parameters of the counting network guidance for a specific text prompt. Although fixed scale parameters can help to control the number of objects to a certain degree, generating the exact number of each object may require tuning the underlying scale parameters.
Second, we found that generating the exact number of more complex objects is much more complicated than generating simple object shapes. The structure of the result image is mostly determined at the early denoising steps, limiting the dividing or merging of objects by counting guidance.
Conclusions
@@
In this paper, we propose counting guidance which is, to our knowledge, the first attempt to guide Stable Diffusion with a counting network to generate the correct number of objects during text-to-image generation. For a single object type, we calculate the gradients of a counting network and refine the estimated noise at every step. For multiple object types, we discuss the semantic information mixing problem and propose attention map guidance to alleviate it. Finally, we obtain masks of each object from the corresponding attention map and calculate the counting network’s gradient of each masked image separately. We demonstrate that our method effectively controls the number of objects with a few limitations. For future work, we will aim to remove the necessarity of scale parameter tuning and create one global framework working for every prompt without additional tuning.

### v3

@@ 4.2.2 Attention Map Guidance
The second row of Fig. 3 demonstrates the effectiveness of our attention map guidance. We find that the attention map for “oranges” focuses solely on oranges, and the attention map for “eggs” focuses solely on eggs, resulting in a correctly synthesized output. Moreover, we observe that high-fidelity object masks are generated from the corresponding attention maps.
4.2.3 Masked Counting Guidance
@@
$M^{b}_{i,j,k}=\begin{cases}1,&\textrm{if}\ i=\operatorname*{argmax}_{i}(M_{i,j,k})\\ 0,&\textrm{otherwise}\end{cases}$ (15)
@@
$L_{count,i}=\left|\frac{Count(\hat{x}_{0,i})-N_{i}}{N_{i}}\right|^{2}.$ (17)
@@
$\epsilon\leftarrow\epsilon+\sum_{i}s_{count,i}\sqrt{1-\bar{\alpha}_{t}}\nabla_{z_{t}}L_{count,i},$ (18)
@@
Stable Diffusion [33]
[…]
Attend-and-Excite [2]
[…]
Ours
[…]
Figure 4: Qualitative comparison for single object type. The first row shows the results of Stable Diffusion [33], the second row shows the results of Attend-and-Excite [2] and the last row shows the results of our method.
5 Experiments
We borrow the state-of-the-art text-to-image generation model Stable Diffusion (v1.4 and v2.1) for our experiments. We use DDIM sampling with 50 steps and set the scale parameter for $L_{max}$ to $s_{max}=0.1$ by default. We create a modified dataset based on the object classes from Attend-and-Excite [2] to evaluate and compare our approach with previous methods. Specifically, we remove the color category and add more animals and objects for a total of 34 object classes. We compare our method with Stable Diffusion [33], Attend-and-Excite [2], and SUR-Adpater [48].
5.1 Quantitative Results
For quantitative comparison, we count the number of given objects using the object detection network Grounding DINO [21]. We create a dataset of 680 prompts using our 34 predefined object classes with counts ranging from 1-20 (e.g., “ten apples”) and measure the normalized MAE (Mean Absolute Error) and RMSE (Root Mean Squared Error). In our evaluation of $s_{count}$, we explored both constant and linearly scheduled approaches. For the constant scenario, we fixed $s_{count}=1$. However, when implementing a linear schedule, we discovered that $s_{count}=\max(0.01,0.2N-1)$ resulted in markedly improved performance. This formulation allows $s_{count}$ to increase incrementally with $N$, providing a more dynamic adjustment compared to the static nature of the constant value (see supplementary materials for detailed hyperparameter analysis).
Tab. 1 presents a detailed quantitative comparison of counting performance. Our method (linear) achieves the best scores for both MAE and RMSE while maintaining comparable or better CLIP similarity to vanilla Stable Diffusion (Tabs. 1(a) and 1(d)). Our method (constant) achieves the second-best score for both MAE and RMSE, demonstrating the effectiveness of our method with fixed $s_{count}$. For the user study, we conducted 330 comparisons on our dataset. In non-tie cases, our method is preferred about 1.9 times more than vanilla Stable Diffusion (Tab. 1(b)).
Despite our method demonstrating superior performance across various metrics, CLIP alone is insufficient to fully reflect image quality, and user studies lack scalability. To address these issues, we incorporate GPT-4V [25] evaluation to further validate the effectiveness of our approach (as shown in Tab. 1(b)). The results indicate that GPT also favors our method over Stable Diffusion, reinforcing the advantages of our strategy.
We also show the effectiveness of our attention map guidance by evaluating text-image and text-text CLIP [30] similarities. We generate 1122 multiple object prompts using our 34 object classes by combining two object classes with a random count for each prompt (e.g., “eight lemons and seventeen onions”). We measure text-image CLIP similarities for all prompts and text-text CLIP similarities for generated captions by BLIP [18] following [2]. We fix the scale parameter to $s_{attention}=1$. Tab. 1(c) presents the quantitative results for both metrics. Attend-and-Excite achieves the best text-image similarity, while our method achieves the best text-text similarity.
5.2 Qualitative Results
@@
Fig. 4 shows a qualitative comparison for the single object type scenario. While Stable Diffusion and Attend-and-Excite fail to generate the right number of objects as specified in the prompt, our method generates the correct number. For the prompt “four tomatoes on the table,” Stable Diffusion generates only three tomatoes without counting guidance. With counting guidance, the tomato at the bottom is successfully divided into two tomatoes, while the rest of the image is consistent with the original result. The text prompt “ten oranges on the table,” causes Stable Diffusion to only generate four oranges compared to our solution that creates the correct amount of ten. The big difference in object count between Stable Diffusion and the target prompt causes large gradients, making our result severely differ from the original.
Our method also works well for more complex categories, such as animals. Considering the prompt “three chicks on the road”, Stable Diffusion and Attend-and-Excite synthesize only two chicks, unlike our method which generates one additional chick while maintaining the other two chicks’ appearance. For the prompt “five rabbits on the yard” Stable Diffusion and Attend-and-Excite generate only four rabbits, while our method generates one more rabbit but fails to maintain the other rabbits’ appearance. That is because of the difference between the background and the rabbit colors. It is hard to generate a white rabbit from a brown yard, so Stable Diffusion with counting guidance changes the overall structure and recreates five rabbits.
@@
Fig. 5 shows a qualitative comparison for multiple object types. For “three lemons and one bread on the table”, Stable Diffusion successfully generates one bread but fails with three lemons, while Attend-and-Excite fails in both cases. With masked counting guidance, our method correctly generates three lemons and one bread. The result shows that the lemon at the bottom is divided into two lemons thanks to masked counting guidance while maintaining the bread’s shape.
@@
For “two onions and two tomatoes on the table”, Stable Diffusion suffers from the semantic information mixing problem and generates red onions instead of tomatoes. Due to our attention map guidance, our method creates realistic tomatoes. As Attend-and-Excite is also based on attention map optimization, it successfully generates realistic tomatoes but fails to generate the exact number of onions.
Baseline Method MAE $\downarrow$ RMSE $\downarrow$ CLIP $\uparrow$
Stable Diffusion Vanilla 0.599 0.746 0.316
Attend-and-Excite 0.601 0.709 0.313
SUR-Adapter 0.903 0.924 0.236
Ours (constant) 0.585 0.696 0.311
Ours (linear) 0.567 0.692 0.315
(a) Counting error and CLIP similarity. Tested with Stable Diffusion.
Baseline Evaluation Tie Vanilla Ours (linear)
Stable Diffusion User study 64.9% 12.1% 23.0%
GPT evaluation 38.5% 26.2% 35.3%
(b) User study and GPT evaluation. Tested with Stable Diffusion.
Baseline Method Text-Image $\uparrow$ Text-Caption $\uparrow$
Stable Diffusion Vanilla 0.324 0.722
Attend-and-Excite 0.330 0.731
SUR-Adpater 0.238 0.563
Ours 0.329 0.732
(c) Effectiveness of attention map guidance. Tested with Stable Diffusion.
Baseline Method MAE $\downarrow$ RMSE $\downarrow$ CLIP $\uparrow$
Stable Diffusion 2 Vanilla 0.473 0.607 0.324
Ours (linear) 0.461 0.593 0.326
(d) Counting error and CLIP similarity. Tested with Stable Diffusion 2.
Table 1: Quantitative results. Evaluated on 680 images.
Failure Cases.
Fig. 6 highlights some failure cases of our method concerning the selection of $s_{count}$. For the prompt “eighteen suitcases”, the vanilla Stable Diffusion generates only four suitcases. Given the large gap between eighteen and four, with $s_{count}=1$, our method adds only one additional suitcase. Increasing $s_{count}$ to 3 results in more suitcases, but it compromises the structure and quality of the image. At $s_{count}=10$, the image becomes significantly distorted. These results emphasize the critical importance of careful hyperparameter selection.
“eighteen suitcases”
[…]
Figure 6: Failure Cases.
6 Limitations
@@
As our results show, our method aids in generating the exact number of each object. However, it is often necessary to tune the scale parameters of the counting network guidance for a specific text prompt (Fig. 6). Although constant or linear scheduling of $s_{count}$ can help to control the number of objects to a certain degree, generating the exact number of each object may require tuning the underlying scale parameters.
7 Conclusions
@@
In this paper, we proposed counting guidance, which, to our knowledge, is the first attempt to guide Stable Diffusion with a counting network to generate the correct number of objects. For a single object type, we calculate the gradient of a counting network and refine the estimated noise at every step. For multiple object types, we discuss the semantic information mixing problem and propose attention map guidance to alleviate it. Finally, we obtain masks of each object from the corresponding attention map and calculate the counting network’s gradient for each masked image separately. We demonstrated that our method effectively controls the number of objects. For future work, we will aim to remove the occasional need for hyperparameter tuning and ensure the framework works more robustly for any prompt.
Acknowledgements
This work was supported in part by Institute of Information & communications Technology Planning & Evaluation (IITP) under the Artificial Intelligence Convergence Innovation Human Resources Development (IITP-2024-RS-2023-00255968) grant funded by the Korea government (MSIT) ITRC and by the MSIT (Ministry of Science and ICT), Korea, under the ITRC (Information Technology Research Center) support program (IITP-2024-2020-0-01461) supervised by the IITP (Institute for Information & communications Technology Planning & Evaluation).
Appendix A Supplementary
This supplementary section provides more information about our experiments, evaluation methods and additional quantitative and qualitative results. We describe in detail how we generate our two evaluation datasets and how we calculate the counting performance of our and previous approaches. We additionally provide more quantitative results to visualize the impact of the choice of our hyperparameters. Finally, we provide a further rich qualitative comparison of our method, Stable Diffusion and Attend-and-Excite to show that our approach outperforms existing ones in various scenarios.
A.1 Dataset
We create two separate datasets for measuring counting loss guidance evaluated by our counting metric and attention loss guidance evaluated by text-image/text-text similarity. The dataset for counting evaluation consists of prompts of a single object with a specific object count. We utilize the 34 object classes from Tab. 2, providing a good balance between simpler to generate objects like fruits and more complex objects like animals. We cover a broad range of object counts ranging from 1-20 per object class to test and compare our method to previous ones. We generate 680 prompts (20 different counts times 34 objects) with the template of the form “{count} {object}” to construct prompts like “one apple”, “three lemons” and “six onions”.
For evaluating our attention loss guidance we use the same 34 objects and build prompts containing two object classes per prompt. Specifically, we form object pairs by combining each object with each other disregarding order and create two prompts per pair with a random count for each object ranging from 1-20. This results in a total of 1122 prompts. We use the template “{count_a} {object_a} and {count_b} {object_b}” yielding examples like “ten cats and five birds”, “nineteen birds and eight lemons” and “five elephants and twelve chicks”.
Table 2: Dataset
Animals
cat, dog, bird, bear, lion, horse, elephant,
monkey, frog, turtle, rabbit, mouse, chick
Objects
backpack, glasses, crown, suitcase, chair,
balloon, bow, car, bowl, bench, clock, apple,
banana, donut, orange, egg, tomato, lemon,
macaron, bread, onion
A.2 Testing Environment
For our experiments, we use PyTorch [27] with a single NVIDIA Tesla V100 32GB GPU. It takes about 12 seconds to generate one image with vanilla Stable Diffusion, while our method takes about 26.9 seconds when using counting guidance for a single object. For two object classes it takes 15 seconds when using attention map guidance only and 37.6 seconds when using both attention map guidance and counting guidance.
A.3 Counting Metric
To calculate our counting metric, we use the state of the art pretrained object detection model Grounding DINO [21] with Swin-T [22] backbone to detect bounding boxes in the generated images. We use the fact that Grounding DINO is able to perform object detection with arbitrary class labels specified as prompts and thus use the objects in the prompt as detection classes. After detection, we count the number of output boxes per object class and compare it with the ground truth count in the prompt. To balance the influence of small and large object counts on the final metric, we additionally normalize our metric by the ground truth object count. Our normalized MAE metric for one object class is given as
$\textit{MAE}=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{\hat{y_{i}}-y_{i}}{y_{i}}\right|,$ (19)
while our normalized RMSE metric is defined as
$\textit{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(\frac{\hat{y_{i}}-y_{i}}{y_{i}}\right)^{2}},$ (20)
where $y_{i}$ is the ground truth object count from the prompt and $\hat{y_{i}}$ is the number of detected bounding boxes in the generated image for the respective class.
A.4 Hyperparameter Analysis
Counting Loss Scale
[…]
Our analysis yielded $s_{count}=\max(0.01,0.2N-1)$, which is a simple increasing function of $N$ that significantly improves performance compared to a constant value.
Attention Loss Scale
Similarly, we visualize the text-text and text-image similarity on our 1122 multi object class dataset for various attention loss scales in Fig. 8. We notice a strong peak of text-text similarity at the value 1 and thus choose our attention loss scale for our experiments as 1.
A.5 Additional Qualitative Results
Fig. 9, Fig. 10 and Fig. 11 show additional results for our counting guidance with various prompts and varying object count for Stable Diffusion, Attend and Excite and ours. Even though we need to tweak our counting guidance scale hyperparameter for some prompts, our counting guidance method consistently creates the correct amount or, when dealing with large count, a similar amount of objects, whereas Stable Diffusion and Attend and Excite fail in many cases. When the object count grows, it becomes more challenging to generate the exact amount, however, our method nevertheless outperforms the other two tested methods.
Fig. 12 visualizes the attention map per object for several prompts for Stable Diffusion and our attention map guidance. We note that our attention maps capture the spatial location of each object more accurately than Stable Diffusion, while reducing the overlap between different objects.
A.6 Template for User Study and GPT Evaluation
User Study
Compare the first and second images provided, and select the one that more closely aligns with the given prompt. Pay particular attention to the object count.
GPT Evaluation Prompt
Compare the first and second images provided, and select the one that more closely aligns with the given prompt. Pay particular attention to the accuracy of the object count. Your selection can be subjective. Your final output score must be either 0 (if the first image is best), 0.5 (’Tie’), or 1 (if the second image is best). You have to give your output in this way (Keep your reasoning concise and short. Give your intermediate thinking step by step.)
$10^{-1}$$10^{0}$$10^{1}$0.70.70.70.70.80.80.80.8counting loss scale $s_{count}$counting error (RMSE)RMSE
[…]
$10^{-1}$$10^{0}$$10^{1}$0.60.60.60.60.70.70.70.7counting loss scale $s_{count}$counting error (MAE)MAE
[…]
55551010101015151515202020200.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.8number of object $N$counting error (MAE)$s_{count}=0.0$$s_{count}=0.1$$s_{count}=0.3$$s_{count}=1.0$$s_{count}=3.0$
[…]
55551010101015151515202020200.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.8number of object $N$counting error (MAE)$s_{count}=0.0$$s_{count}=0.1$$s_{count}=0.3$$s_{count}=1.0$$s_{count}=3.0$
[…]
Figure 7: Hyperparameter study. Evaluated on 680 images.
$10^{-1}$$10^{0}$$10^{0.7}$0.320.320.320.320.330.330.330.33attention loss scalesimilaritytext-image
$10^{-1}$$10^{0}$$10^{0.7}$0.720.720.720.720.740.740.740.74attention loss scalesimilaritytext-text
Figure 8: Effect of attention loss scale on the text-image and text-text CLIP similarity. Evaluated on our 1122 two object prompt dataset.
Stable Diffusion
[…]
Attend-and-Excite
[…]
Ours
[…]
Figure 9: Additional qualitative results (1)
Stable Diffusion
[…]
Attend-and-Excite
[…]
Ours
[…]
Figure 10: Additional qualitative results (2)
Stable Diffusion
[…]
Attend-and-Excite
[…]
Ours
[…]
Figure 11: Additional qualitative results (3)
Stable Diffusion
“apples and donuts on the table”
[…]
“strawberries and eggs on the table”
[…]
Ours
“apples and donuts on the table”
[…]
“strawberries and eggs on the table”
[…]
Figure 12: Additional qualitative results (4)